DX Reconstructor Measurements

16:43
  • Looking for Source
17:28
  • Moving stage to get rid of dust motes...AGAIN!
  • Cube stage to 1.4,
  • Cub rotator: 203.708
  • Baysides -75.5 -49.4 48.492
  • Hexapod -1.6 -0.947 4.4 25.2 22 0
17:30
  • Taking new flat with reconstructor: /local/aomeas/adsec_calib/M2C/KL_v16/RECs/Rec_20170612_230117.fits, yesterday's last
  • Flat: 20170613_400_flat.sav
  • Same disturb as yesterday's last:20170612_ 225127
  • Intmat, 48 iter: 173519
  • Recon: 175001
17:53
  • Testing with turbulence.
  • First, to find the best focus on LUCI2
  • Best focus 50.1
18:38
  • Testing with turbulence.
  • 0.8" /local/aomeas/adsec_calib/M2C/KL_v16/disturb/CmdDisturb_20170613_184455.fits
  • No GOPT
  • Manual gains 0.60 0.40 0.30
  • Datasets: 184834, 184852, 184909, 185019
  • NCPA on
  • GOPT 0.93
  • Rotator angle 210, dividing amplitudes by 2 as the system is double pass. Pa_beta
  • Datasets: 185322, 185517, 185535, 185552
18:56
  • 1.0" dist_flao1_ts4_atm_s1.0_L040.0_v15.0_ovfreq1000.00_sd3892.fits, adding sinus mode: 20170613_185854
  • Gains 0.65 0.40 0.35
  • Datasets: 190135, 190153, 190215, 190240
  • NCPA on
  • GOPT 0.85
  • Datasets: 190347, 190407, 190432, 190454
19:06
  • 1.2" dist_flao1_ts4_atm_s1.2_L040.0_v15.0_ovfreq1000.00_sd3892.fits, adding sinus mode: 20170613_190727
  • Gains: 0.80 0.50 0.40
  • Datasets: 190910, 190928, 191011, 191031
  • NCPA on
  • Datasets: Cannot apply NCPA, one side became dark and the edge rolled.
20:05
  • Trying the 10 iteration ff matrix
  • Having trouble getting light in the bottom edge. Saving new flat and continuing to try. 20170613_400_flat_2.sav
  • Trying 300 hz, 12 l/d
20:50
  • Many issues with the wfs. I had to restart. There was an error to the effect that a slot in the shared MASTERDIAGNOSTIC buffer was unavailable and I couldn't take darks on CCD39
21:00
  • System recovered and lamp still works.
  • New flat at 300hz and 12 l/d: 20170613_400_flat_3.sav
  • The latest reconstructor is not terribly stable with the 10 iteration ff.
  • flat: 20170613_400_flat_4.sav
  • New 4 iteration intmat: 211434
  • Recon: 211604
21:19
  • 0.8" disturb 184455
  • Gains: 0.60 0.40 0.30
  • Datasets: 212305, 212416, 212437, 212459
  • NCPA on
  • GOPT 0.9
  • 212919, 212938, 213004, 213027
21:30
  • 1" disturb 185854
  • gain 0.65 0.40 0.35
  • Datasets: 213223, 213240, 213301, 213320
  • NCPA on
  • GOPT 0.8
  • 213358, 213419, 213439, 213524
21:36
  • 1.2" disturb 190727
  • Gains 0.80 0.50 0.40
  • Loop unstable at any ho2 gain.
21:39
  • Trying the old FF matrix with a 2015 reconstructor that works particularly well:
  • Saved the 5 iteration ff matrix in/home/aoeng/act_add/calib/adsec/current/data_5, the 10 iteration in .../data_10
  • The old ff matrix (in .../data_old) has been copied to .../data
  • Adsec started in an unknown state
  • HBS may have been off for the last couple of datasets.
  • stopped IDL controller, fast diagnostic then back on in reverse order,
  • attempted recover fail, it didn't work, restarted adsec arbitrator now the system is in power off.
  • The AdSec control gui froze. Started a new one and began power on.
  • power on froze
  • the copy from data_old to data was bad. I should have deleted data first.
  • Now it's powering on.
22:10
  • Went into safe skip and failed to set the flat, that's a new one.
  • Trying: idl> print, fsm_set_flat(/no_flat), then setting a flat other than default.
  • Still goes into safe skip!
  • Cannot set shell with this FF matrix!!!!
22:19
  • Going back to the 10 iteration ff matrix to try and re-take the 1.2" data, since HBS was off.
  • Actuator 557 is bad, reading 140 microns. Put in the no position list.
  • Using 1.2" disturb the instability might have been from that bad actuator.
  • Gains 0.80 0.50 0.40
  • NO NCPA
  • Datasets: 224326, 224342, 224359, 224421
  • NCPA on
  • GOPT 0.72
  • Still goes into safe skip. Just in case, I'll take a 48 iter intmat.
  • Intmat: 224815, Taken at 990 Hz, ARRGH.
  • intmat: 230329
  • Recon: 231742
23:21
  • 1.2" again,
  • Gain 0.80 0.50 0.40.
  • No NCPA
  • 232202, 232222, 232240, 232258
  • NCPA on
  • GOPT 0.7
  • 232416, 232435, 232453, 232517
23:25
  • Ramping down. getting ready for handover.
Image of pupil after setting the flat, the images are stretched, the pupils are actually fairly round:

Image of pupil with the loop running, 0.8":